OpenStudy (anonymous):

8 times the square root of 63 x to the fifth power

OpenStudy (anonymous):

24 √7 x^5

OpenStudy (anonymous):

Assume that the expression is 8√(63x^5). Since 63x^5 contains the perfect square factors of 9 and x^4. 8√(63x^5) = 8√(9x^4 * 7x) = 8 * 3x^2√(7x) = 24x^2√(7x)

OpenStudy (anonymous):